Use Cramer's rule to solve the unknown of the system: x + 2y =2 - z,3x - 6y = 2 - 2x,2x =8 + z


i.x =2, y=1/3, z=-2

ii.x =-3,y =2, z=-1/2

iii.x =4,y =-1/2, z =-3

iv.x =3, y =1/2, z=-2




1
Expert's answer
2021-08-26T01:06:53-0400

"\displaystyle 
x + 2y +z=2
\\5x-6y=2
\\2x-z=8
\\\text{We can write this in matrix form}
\\\begin{pmatrix}
1&2&1\\
5&-6&0\\
2&0&-1
\end{pmatrix}\begin{pmatrix}
x\\y\\z
\end{pmatrix}=\begin{pmatrix}
2\\2\\8
\end{pmatrix}
\\\text{Next,we use the Cramer's rule to to solve the system of equations}
\\detA = 1\begin{vmatrix}
5&-6\\
2&0
\end{vmatrix}-1\begin{vmatrix}
1&2\\
5&-6
\end{vmatrix}=28
\\\delta_x=\begin{vmatrix}
2&2&1\\
2&-6&0\\
8&0&-1
\end{vmatrix}=64
\\\delta_y=\begin{vmatrix}
1&2&1\\
5&2&0\\
2&8&-1
\end{vmatrix}=44
\\\delta_z=\begin{vmatrix}
1&2&2\\
5&-6&2\\
2&0&8
\end{vmatrix}=-96
\\\text{Therefore using Kramer's rule we have that, $\overline{x}=\frac{64}{28}=2.29$}
\\\overline{y} = \frac{44}{28}=1.57
\\\overline{z} = \frac{-96}{28} =-3.42
\\\text{Hence, the values of x,y and z are 2, 2 and -3 approximately}"